Ticket Number: HEA-NOC/20070910-5 Ticket Status: UPDATE
Ticket Type: unscheduled Resolver: HEA-NOC
Ticket Opened: 20070910 16:17 UTC+1 Problem Start: 20070910 14:14 UTC+1
Ticket Update: 20070911 17:14 UTC+1 Problem End:
Site/line: TCD Tallaght Hospital, UCD
The UPC fibre between UCD, IT Tallaght, Tallaght Hospital has gone down.
Traffic from UCD to the HEAnet backbone normally flows over this link, and
has rerouted normally.
Traffic from TCD to Tallaght Hospital is currently not flowing.
20070910 16:22 BB Explained problem to NTL/UPC, and requested diagnosis
20070910 16:23 BB added TCD's Tallaght Hospital to the affected list
20070910 16:43 BB UPC report there are no other faults on parallel
circuits, and requested we check power/patching on the affected fibre
20070911 09:55 BB LT has checked our equipment on site, and local
patching. All seems normal - asking UPC to carry out tests on the fibre.
20070911 13:29 LT UPC engineer rang for further info. He will run an OTDR
trace on the link to troubleshoot the problem, from Tallaght Hospital end.
He expects to be at Tallaght Hospital within the hour.
20070911 15:01 BB UCD plan to change the OSPF costs on their links to make
the use of the link via the TCD pop the primary.
20070911 15:42 LT UPC engineer is on-site at Tallaght Hospital. He can't
see any light on the link when he attaches a light meter to it. I have
shut down the switch interfaces at EsatX end so that he can run an OTDR
trace on the link.
20070911 15:56 LT UPC engineer has identified a possible fibre break
approximately 610 metres from Tallaght Hospital. He is carrying out a test
on the fibre leg to IT Tallaght to verify that his OTDR kit is calibrated
correctly - if these tests produce expected results, then it supports the
findings on the leg to EsatX in which case he will report back to UPC so
that a splicing crew can be organised.
20070911 16:26 LT Further tests support the results suggesting a fibre
break near to Tallaght Hospital. UPC are currently looking at scheduling a
crew to investigate.
20070911 17:02 BB UCD and HEAnet-NOC have reversed the OSPF costs to
prevent potential traffic flow flaps
20070911 17:07 LT UPC expect to have a splice crew look at the break this
evening. I have brought the switch interfaces at EsatX back up again, at
the suggestion of UPC engineer, so that they can match transmit and
receive when re-joining the fibre strands.
Based on current evidence, the break is possibly in a UPC chamber just
outside the hospital grounds. If it does prove to be in a chamber, then we
would hope that the fibre will be repaired tonight and usable by tomorrow
morning. Otherwise, if the break is between chambers, resolution may be
more complex and therefore more time consuming.
This ticket can be monitored at http://www.hea.net/tickets/20070910-5
HEAnet Limited. Registered in Ireland, No. 275301.